OpenStudy (anonymous):

Is the sequence 6,12,24,48... geometric? If so, identify the common ratio

OpenStudy (anonymous):

A geometric sequence is a sequence (finite or infinite) of real numbers for which each term is the previous term multiplied by a constant (called the common ratio). For example, starting with 3 and using a common ratio of 2 we get the finite geometric sequence: 3, 6, 12, 24, 48; and also the infinite sequence...

OpenStudy (anonymous):

Does that help if not I will show you something else

OpenStudy (confluxepic):

Yes. @~ @! @_ @, @; @confluxepic

OpenStudy (anonymous):

yes;2 6*2=12 12*2=24 24*2=48